OpenStudy (1018):

differentiate e^(x+y) = ln(x+y)+3x+3y+3

OpenStudy (sidsiddhartha):

use chain rule in both sides

OpenStudy (sidsiddhartha):

first diffrentiate LHS can u do it?

OpenStudy (1018):

Ok this what I got. Please see if it's correct. I'll do the left first. :)

OpenStudy (sidsiddhartha):

yeah go ahead

OpenStudy (1018):

\[e ^{x+y} y\prime\]

OpenStudy (1018):

Is this correct?

OpenStudy (zzr0ck3r):

what are you differentiating with respect to?

OpenStudy (zzr0ck3r):

is this implicite?

OpenStudy (sidsiddhartha):

almost theres a slight mistake look-- \[\frac{ d }{ dx }e^{x+y}=e^{x+y}*\frac{ d }{ dx}(x+y) =e^{x+y}*(1+y^{'})\] ok?

OpenStudy (sidsiddhartha):

got it? you missed that x

OpenStudy (1018):

yes implicate

OpenStudy (1018):

oh yeah, crap I forgot that one. Thanks. :)

OpenStudy (sidsiddhartha):

now try RHS

OpenStudy (1018):

\[\frac{ y\prime\ }{ x+y } +3 + 3y\prime\\] How's this?

OpenStudy (1018):

Oh pellet wait wth

OpenStudy (1018):

lol pellet i cant say sh*t here lolol

OpenStudy (1018):

\[\frac{ y\prime\ }{ x+y } +3 + 3y\prime\\]

OpenStudy (sidsiddhartha):

uhh yout latex is'nt working

OpenStudy (1018):

Crap. can you try copying it to yours? maybe it'll show up.

OpenStudy (sidsiddhartha):

ok i'm doing it ,check it with yours

OpenStudy (sidsiddhartha):

\[\frac{ d }{ dx }[\ln(x+y)+3x+3y+3]=\frac{ 1 }{ x+y }\frac{ d }{ dx }(x+y)+3+3y^{'}+0\\=\frac{ 1 }{ x+y }*(1+y^{'})+3+3y^{'}\] ok?

OpenStudy (sidsiddhartha):

@1018

OpenStudy (1018):

i forgot the 1 again. tsk. wait, cant it be 1+y(prime) / x +y ?

OpenStudy (1018):

cant i put the 1+yprime as the numerator?

OpenStudy (sidsiddhartha):

no you cant because u're diffrentiating (x+y) alone after differentiating it with log

OpenStudy (1018):

oh. ok. cause in my notes, it says the formula is du/u is it wrong?

OpenStudy (sidsiddhartha):

itd like \[\frac{ d }{ dx }\ln(x^2+x^3)=\frac{ 1 }{ x^2+x^3 }*(2x+3x^2)\] like this

OpenStudy (1018):

yeah that, but in mine it's du/u. so is it wrong to be \[\frac{ 2x + 3x^2 }{ x^2 + x^3}\]

OpenStudy (sidsiddhartha):

no its not wrong

OpenStudy (1018):

oh. ok thanks for clarifying. so now, i got \[e ^{x+y} 1+y\prime\ = \frac{ 1+y\prime }{ x + y } + 3 + 3y\]

OpenStudy (1018):

correction. that's 3yprime

OpenStudy (sidsiddhartha):

yes thats correct :)

OpenStudy (1018):

then what's next :)

OpenStudy (sidsiddhartha):

then take things with y(prime) to a side take take y(prime) common

OpenStudy (sidsiddhartha):

like this-- \[e^{x+y}+e^{x+y}*y^{'}=\frac{ 1 }{ x+y }+\frac{ y^{'} }{ x+y }+3+3y^{'}\\y^{'}[e^{x+y}-\frac{ 1 }{ x+y }-3]=\frac{ 1 }{ x+y }+3-e^{x+y}\]

OpenStudy (sidsiddhartha):

got it?

OpenStudy (1018):

wait, haha why are there two e in the lhs?

OpenStudy (sidsiddhartha):

because we got \[LHS=e^{x+y}(1+y^{'})=e^{x+y}+e^{x+y}*y^{'}\\just ~multiplied~them\]

OpenStudy (1018):

oh yeah yeah. crap man i need to it i'm losing focus. thanks for the help. the final answer for this is -1, right?

OpenStudy (1018):

*eat

OpenStudy (sidsiddhartha):

i dont know you can calculate it easy algebra only :)

OpenStudy (1018):

no our professor told us that it should be -1 lol

OpenStudy (1018):

ok ok thanks appreciate your help!

OpenStudy (sidsiddhartha):

yes it will be -1 just calculate it u'll get the answer

OpenStudy (sidsiddhartha):

yw!!!